OpenStudy (anonymous):

Simplify the following complex fraction. (PROBLEM WILL BE POSTED ON QUESTION) A.16/35 B.7/20 C.5/7 D.7/5

OpenStudy (anonymous):

\[\frac{ 0.4 }{ 0.5\times \frac{ 7 }{ 4}}\]

OpenStudy (anonymous):

@ganeshie8 @chosenmatt @Marki @TheAsker2002

OpenStudy (chosenmatt):

what do you think?

OpenStudy (anonymous):

um B?

OpenStudy (chosenmatt):

yup good job!!!

OpenStudy (anonymous):

Wrong one

OpenStudy (anonymous):

its A

OpenStudy (chosenmatt):

oh :(

OpenStudy (anonymous):

:P